OpenStudy (adamaero):

2 instructors grade on a the curve, that is, give a fixed percentage of each various letter grades to each of their classes. Their curves are as follows: Grade Instructor A Instructor B A 10% 15% B 15% 15% C 45% 30% D 15% 20% F 15% 20% Question: If a random student came to you and said that his object was to enroll in the class in which he could expect the higher grade point average, which instructor would you recommend?

OpenStudy (adamaero):

@tkhunny @nincompoop

OpenStudy (tkhunny):

As with all such things, Average = Mean = Expected Value = (Possible Values)*(P(Possible Values)) Go!

OpenStudy (adamaero):

Can you explain a little more please? 5*(0.1+0.15+0.45+0.15+0.15)

OpenStudy (amistre64):

spose we define the grades as 0,1,2,3,4 this way we can have some numerical determinations

OpenStudy (amistre64):

E(X) = sum of x p(x)

OpenStudy (amistre64):

consider it this way; out of 100 grades a score of 0 is expected to happen 15 times out of 100 a score of 1 is expected to happen 15 times out of 100 a score of 2 is expected to happen 45 times out of 100 a score of 3 is expected to happen 15 times out of 100 a score of 4 is expected to happen 10 times out of 100 how do we determine an average? (sum of grades)/#of grades 15 0s is just 15(0) 15 1s is just 15(1) 45 2s is just 45(2) 15 3s is just 15(3) 10 4s is just 10(4) \[\frac{0(15)+1(15)+2(45)+3(15)+4(10)}{100}\] \[0\frac{15}{100}+1\frac{15}{100}+2\frac{45}{100}+3\frac{15}{100}+4\frac{10}{100}\] \[0P(0)+1P(1)+2P(2)+3P(3)+4P(4)\] \[\sum_{k=1}^{n}x_kP(x_k)\]

OpenStudy (adamaero):

ohhh, I got it now! =D assign grade F with zero & go up from there which makes a lot of sense when a 4.0 GPA is an A

OpenStudy (amistre64):

:) yep